# SS Linear Fusion Model: Hyperspectral Imaging Efficient Spatial and Spectral Linear Model with Bidirectional Feature Learning

### v1

@@ Abstract
Classifying hyperspectral images (HSIs) is a complex task in remote sensing due to the high-dimensional nature and volume of data involved. To address these challenges, we propose the Spectral-Spatial Linear (SS Linear) Model, a novel framework that significantly reduces data volume while enhancing classification accuracy. Our model employs a bidirectional reversed convolutional neural network (CNN) to efficiently extract spectral features, complemented by a specialized block for spatial feature analysis. This hybrid approach leverages the operational efficiency of CNNs and incorporates dynamic feature extraction inspired by attention mechanisms, optimizing performance without the high computational demands typically associated with transformer-based models. The SS Linear Model is designed to process hyperspectral data bidirectionally, achieving notable classification and efficiency improvements by fusing spectral and spatial features effectively. This approach yields superior classification accuracy compared to existing benchmarks while maintaining computational efficiency, making it suitable for resource-constrained environments. We validate the SS Linear Model on three widely recognized datasets, Houston 2013, Indian Pines, and Pavia University, demonstrating its ability to outperform current state-of-the-art models in HSI classification and efficiency. This work highlights the innovative methodology of the SS Linear Model and its practical benefits for remote sensing applications, where both data efficiency and classification accuracy are critical. For further details, please refer to our code repository on GitHub: HSILinearModel.
@@
Hyperspectral image,Bidirectional networks, Feature extraction, Classification, Computing efficiency, Spectral and spatial linear fusion.
@@ I Introduction
In response, we introduce the Spectral-Spatial Linear (SS Linear) model, a novel architecture designed to enhance HSI classification through a fusion of CNN and attention-based strategies. Our model employs a bidirectional network to efficiently process hyperspectral data, integrating both spectral and spatial information into a unified representation. By leveraging CNNs for spatial feature extraction alongside Mamba-inspired mechanisms for reduced computational load, the SS Linear Model achieves robust classification performance with lower memory and computational costs. This architecture supports a comprehensive understanding of both local and global patterns in HSI data, effectively addressing the high-dimensional nature of hyperspectral images without the intensive computational overhead associated with transformers.
The SS Linear Model was rigorously evaluated on three prominent hyperspectral datasets—Houston 2013, Indian Pines, and Pavia University—demonstrating superior classification performance and computational efficiency compared to current state-of-the-art models. By reducing GPU memory usage, CPU load, and inference time, the SS Linear Model represents a promising advancement in HSI analysis, particularly for dense prediction tasks that benefit from efficient processing pipelines.
@@
  • We introduce the SS Linear Model, a novel bidirectional framework for HSI classification that integrates spatial and spectral feature extraction. This model uniquely balances computational efficiency and classification accuracy, setting a new standard for HSI classification architectures.
  • The SS Linear Model achieves computational advantages over existing models, such as, RNN, CNN-based architectures, and ViT-based. This design minimizes memory and processing requirements, facilitating scalable classification suitable for practical remote sensing applications.
  • We validate the performance of the SS linear model through extensive experiments on three major datasets: Houston 2013, Indian Pines, and Pavia University. Our results consistently outperform vision transformer benchmarks, establishing the SS Linear Model as a high-performance, efficient option for hyperspectral image classification.
The rest of this paper is structured as follows. Section 2 reviews related work, situating our approach within the current landscape of HSI classification research. Section 3 details the methodology of the SS linear model, while Section 4 presents experimental results and analyzes. Finally, Section 5 concludes with insight into potential future research directions.
@@ II-C The Emergence of State Space Models: Mamba in Computer Vision Applications
State Space Models (SSMs) have recently gained traction in vision applications as a scalable alternative to transformers. The Mamba model [17], a prominent example of an SSM-based approach, has demonstrated linear scalability and effective handling of long-range dependencies, making it suitable for tasks that require efficient processing across large data volumes. By replacing attention mechanisms with a computationally efficient backbone, Mamba offers a promising solution to the high resource demands of transformer models.
@@ III Methodology
This section presents the proposed SS Linear Model methodology, designed specifically for hyperspectral image (HSI) classification. The SS Linear Model aims to address two critical challenges in HSI analysis: (1) high computational demands due to the large volume of spectral and spatial data and (2) maintaining or improving classification accuracy. By leveraging a unique bidirectional processing framework, our model effectively reduces the computational resources required for HSI classification while simultaneously enhancing performance.
The SS Linear Model is built upon core principles of spectral-spatial feature extraction, guided by linear transformation parameters $A$ and $B$, which enable efficient forward and backward processing along the spectral dimension. The model’s architecture integrates a bidirectional Bi-Networks block for spectral analysis and a spatial processing block for spatial context capture. Together, these components create a comprehensive representation of hyperspectral data, optimized for efficient processing and improved classification.
The following sections describe the mathematical foundations, architectural components, and computational benefits of the SS Linear Model in detail.
III-A SS Linear Model Model Preliminaries
The SS Linear Model is built upon bidirectional spectral processing principles, specifically tailored for hyperspectral image (HSI) classification. The model leverages two transformation matrices,$A$ and $B$, designed to capture forward and backward dependencies across the spectral dimension. This bidirectional approach enables the model to synthesize spectral information both sequentially and retrospectively, thus creating a richer feature representation.
@@ III-A1 Bidirectional Processing Framework
In the SS Linear Model, hyperspectral data $x(t)$ at each spectral band $t$, is projected into a hidden state representation $h(t)$ that integrates both forward and backward information. The forward and backward state transitions are modeled as follows:
@@
This bidirectional processing not only captures sequential dependencies, but also improves the discriminative power of the extracted features. By integrating forward and backward information, the SS Linear Model effectively encapsulates the spectral structure in hyperspectral data, which is critical for accurate classification.
@@ III-B Proposed Method Overview and Architecture Description
The proposed SS Linear Model architecture (illustrated in Figure 1) is designed to efficiently handle the high-dimensional data of hyperspectral images by dividing the processing into distinct blocks that target spectral and spatial information separately. This architecture consists of four main components: 1. Input Patch Preparation; 2. Hyperspectral Bi-Networks Block (for bidirectional spectral processing); 3. Spatial Feature Processing Block; 4. Classifier Block.
@@
Figure 1: The architectural overview of the Proposed SS Linear Model model. The framework consists of four main components: (A) A hyperspectral image patch with dimensions $p\times p\times\text{CH}$; (B) The Hyperspectral Bi-networks Block; (C) The Spatial Processing Block; (D) The classifier. The process begins by extracting patches that serve as input to the SS Linear Model block. This block includes a spatial processing stage that precedes the unique forward and backward operations, and the concatenation between the Bi-networks output and the spatial feature process is the input of the classifier block.
@@ III-B2 Hyperspectral BiNetwork Block
The hyperspectral bi-network block is central to the architecture of the SS Linear Model. It processes the spectral information bidirectionally, utilizing forward and backward transformations to capture spectral dependencies in both directions. The following steps outline the processing within the Bi-Networks Block:
@@ III-B4 Classifier Block
III-B5 Algorithm for SS Linear Model Operation
The SS Linear Model efficiently processes hyperspectral image patches by integrating bidirectional spectral feature extraction and spatial processing to achieve accurate classification. The main steps in this operation involve Input Normalization, Forward and Backward Projections,Bidirectional Spectral Processing,Spatial Feature Processing, and Feature Concatenation and Classification.
Algorithm 1 SS Linear Model Operation
@@ III-B5 Algorithm for SS Linear Model Operation
5:Project $x_{\text{norm}}$ to $x_{\text{proj}}$ and $z_{\text{proj}}$ using linear transformations
@@ III-C Architectural Efficiency Analysis
The SS Linear Model is designed to reduce computational demands while maintaining or improving classification accuracy. This section provides a quantitative analysis of the model’s efficiency, highlighting its advantages in terms of FLOPs, runtime, and overall computational complexity.
@@ III-C1 Floating Point Operations (FLOPs) Reduction
FLOPs provide a measure of the computational load associated with performing a single forward pass in a model. In hyperspectral image classification, conventional CNNs and transformers often involve high FLOPs due to extensive convolutional or self-attention operations across multiple spectral bands and spatial dimensions. The SS Linear Model achieves a substantial reduction in FLOPs by adopting a bidirectional spectral processing approach without extensive kernel operations. Instead of using large convolutional filters or complex attention mechanisms, the SS Linear Model leverages simpler 1D convolutions along with the efficient transformation matrices $A$ and $B$ resulting in lower computational complexity. To summarize, for a hyperspectral image with dimensions $Batch\times H\times W\times C$, where:
@@
The SS Linear Model’s bidirectional approach enables it to reduce FLOPs by approximately 40% compared to conventional CNNs, assuming a kernel size $k>3$. This reduction is even more pronounced when compared to transformers, where the self-attention mechanism has quadratic complexity in the spectral dimension $CH$,
@@ III-C2 Runtime Efficiency
In practical runtime evaluations, the SS Linear Model demonstrates a significant improvement in inference time over transformer-based architectures. This efficiency is primarily attributed to two architectural choices:
@@
Reduced Memory Footprint: The reduced number of parameters in the SS Linear Model lowers memory usage, allowing it to perform faster, particularly on GPU and edge devices with limited memory resources.
When tested under identical hardware conditions, the SS Linear Model achieved a 30% faster inference time compared to transformer-based models, demonstrating its practical advantage for real-time applications.
@@ III-C3 Complexity Reduction
The bidirectional processing framework used in the SS Linear Model further reduces computational complexity by consolidating forward and backward spectral processing into a single framework. This approach contrasts with traditional models that handle spectral and spatial processing in isolation, often leading to redundant computations.
The SS Linear Model achieves efficiency through the following approximations:
@@
    SS Linear Model: Parameters: $O(CH+HW)$, FLOPs: $O(Batch.H.W\cdot CH)$.
The significantly lower FLOPs and parameter count in the SS Linear Model contribute to a leaner architecture, suitable for deployment in scenarios where computational resources are constrained.
@@ IV Experiments
This section presents a comprehensive evaluation of the proposed SS Linear Model, including dataset descriptions, experimental setup, benchmark comparisons, and in-depth analysis of classification performance. Our experiments examine model efficacy across several hyperspectral datasets, benchmarking SS Linear Model against state-of-the-art methods, and include ablation studies to validate its structural advantages.
@@ IV-A Datasets Description
Our experiments are conducted on three prominent hyperspectral datasets—Houston 2013, Indian Pines, and University of Pavia—each offering unique characteristics that make them ideal for evaluating the SS Linear Model’s performance across varied spectral and spatial complexities. The diversity of these datasets provides a comprehensive test of the model’s adaptability and robustness in hyperspectral image classification.
@@ 9 Shadows 103 923 1026
Each dataset’s specific spectral and spatial properties enable a well-rounded evaluation of the SS Linear Model, examining its robustness across complex, heterogeneous, and mixed-land-cover environments.
@@ IV-B Experimental Setup
To comprehensively assess the SS Linear Model’s classification effectiveness, we designed a robust experimental setup, encompassing evaluation metrics, benchmark comparisons, and specific implementation details. This setup provides a clear framework for evaluating both the accuracy and computational efficiency of the proposed model.
@@ IV-B1 Evaluation Metrics
The performance of the SS Linear Model is quantitatively assessed using a suite of standard metrics. Specifically, we employ Overall Accuracy ($OA$), which measures the general precision of the model; and the Kappa coefficient ($\kappa$), a statistical measure that accounts for chance agreement in classification tasks. Together, these metrics enable a robust evaluation of model performance.
@@ IV-B2 Benchmark Comparisons
We benchmark the SS Linear Model against a set of widely recognized hyperspectral image classification models, covering various deep learning architectures. These benchmarks encompass:
@@
To ensure consistency, comparative experiments are conducted with a standardized patch size of 15 and a 10% training data ratio. Additionally, we evaluate SS Linear Model with a smaller patch size of 3 to demonstrate its ability to maintain high accuracy with reduced computational resources, a key advantage in resource-limited environments.
@@ IV-B3 Implementation Details
The SS Linear Model is implemented using the PyTorch platform and trained on the school server based on GPU for enhanced computational performance. Optimization and Training: We use the Adam optimizer with a mini-batch size of 32 and a learning rate of 5e-4, with training set to 100 epochs across all datasets to balance training efficiency and model performance. Loss Function: Cross-entropy loss is chosen due to its effectiveness in multi-class classification tasks, particularly for hyperspectral data where class imbalances often require robust loss formulations. Data Augmentation: To improve model robustness, we apply geometric transformations, including rotations at 45°, 90°, and 135°, as well as vertical and horizontal flips, to the training data. This augmentation strategy enriches the training dataset by introducing additional spatial variability, enhancing the model’s ability to generalize across different spectral-spatial patterns. This experimental setup ensures that the SS Linear Model is rigorously tested under conditions that reflect both practical application scenarios and optimal evaluation standards.

### v2

# Hyperspectral Images Efficient Spatial and Spectral non-Linear Model with Bidirectional Feature Learning

@@ Abstract
Classifying hyperspectral images (HSIs) is a complex task in remote sensing due to the high-dimensional nature and volume of data involved. To address these challenges, we propose the Spectral-Spatial non-Linear Model, a novel framework that significantly reduces data volume while enhancing classification accuracy. Our model employs a bidirectional reversed convolutional neural network (CNN) to efficiently extract spectral features, complemented by a specialized block for spatial feature analysis. This hybrid approach leverages the operational efficiency of CNNs and incorporates dynamic feature extraction inspired by attention mechanisms, optimizing performance without the high computational demands typically associated with transformer-based models. The SS non-Linear Model is designed to process hyperspectral data bidirectionally, achieving notable classification and efficiency improvements by fusing spectral and spatial features effectively. This approach yields superior classification accuracy compared to existing benchmarks while maintaining computational efficiency, making it suitable for resource-constrained environments. We validate the SS non-Linear Model on three widely recognized datasets, Houston 2013, Indian Pines, and Pavia University, demonstrating its ability to outperform current state-of-the-art models in HSI classification and efficiency. This work highlights the innovative methodology of the SS non-Linear Model and its practical benefits for remote sensing applications, where both data efficiency and classification accuracy are critical. For further details, please refer to our code repository on GitHub: HSILinearModel.
@@
Hyperspectral image,Bidirectional networks, Feature extraction, Classification, Computing efficiency, Spectral and spatial non-linear fusion.
@@ I Introduction
In response, we introduce the Spectral-Spatial non-linear (SS non-linear) model, a novel architecture designed to enhance HSI classification through a fusion of CNN and attention-based strategies. Our model employs a bidirectional network to efficiently process hyperspectral data, integrating both spectral and spatial information into a unified representation. By leveraging CNNs for spatial feature extraction alongside Mamba-inspired mechanisms for reduced computational load, the SS non-linear Model achieves robust classification performance with lower memory and computational costs. This architecture supports a comprehensive understanding of both local and global patterns in HSI data, effectively addressing the high-dimensional nature of hyperspectral images without the intensive computational overhead associated with transformers.
The SS non-linear Model was rigorously evaluated on three prominent hyperspectral datasets—Houston 2013, Indian Pines, and Pavia University—demonstrating superior classification performance and computational efficiency compared to current state-of-the-art models. By reducing GPU memory usage, CPU load, and inference time, the SS non-linear Model represents a promising advancement in HSI analysis, particularly for dense prediction tasks that benefit from efficient processing pipelines.
@@
  • We introduce the SS non-linear Model, a novel bidirectional framework for HSI classification that integrates spatial and spectral feature extraction. This model uniquely balances computational efficiency and classification accuracy, setting a new standard for HSI classification architectures.
  • The SS non-linear Model achieves computational advantages over existing models, such as, RNN, CNN-based architectures, and ViT-based. This design minimizes memory and processing requirements, facilitating scalable classification suitable for practical remote sensing applications.
  • We validate the performance of the SS non-linear model through extensive experiments on three major datasets: Houston 2013, Indian Pines, and Pavia University. Our results consistently outperform vision transformer benchmarks, establishing the SS non-linear Model as a high-performance, efficient option for hyperspectral image classification.
The rest of this paper is structured as follows. Section 2 reviews related work, situating our approach within the current landscape of HSI classification research. Section 3 details the methodology of the SS non-linear model, while Section 4 presents experimental results and analyzes. Finally, Section 5 concludes with insight into potential future research directions.
@@ II-C The Emergence of State Space Models: Mamba in Computer Vision Applications
State Space Models (SSMs) have recently gained traction in vision applications as a scalable alternative to transformers. The Mamba model [17], a prominent example of an SSM-based approach, has demonstrated non-linear scalability and effective handling of long-range dependencies, making it suitable for tasks that require efficient processing across large data volumes. By replacing attention mechanisms with a computationally efficient backbone, Mamba offers a promising solution to the high resource demands of transformer models.
@@ III Methodology
This section presents the proposed SS non-linear Model methodology, designed specifically for hyperspectral image (HSI) classification. The SS non-linear Model aims to address two critical challenges in HSI analysis: (1) high computational demands due to the large volume of spectral and spatial data and (2) maintaining or improving classification accuracy. By leveraging a unique bidirectional processing framework, our model effectively reduces the computational resources required for HSI classification while simultaneously enhancing performance.
The SS non-linear Model is built upon core principles of spectral-spatial feature extraction, guided by non-linear transformation parameters $A$ and $B$, which enable efficient forward and backward processing along the spectral dimension. The model’s architecture integrates a bidirectional Bi-Networks block for spectral analysis and a spatial processing block for spatial context capture. Together, these components create a comprehensive representation of hyperspectral data, optimized for efficient processing and improved classification.
The following sections describe the mathematical foundations, architectural components, and computational benefits of the SS non-linear Model in detail.
III-A SS non-Linear Model Model Preliminaries
The SS non-linear Model is built upon bidirectional spectral processing principles, specifically tailored for hyperspectral image (HSI) classification. The model leverages two transformation matrices,$A$ and $B$, designed to capture forward and backward dependencies across the spectral dimension. This bidirectional approach enables the model to synthesize spectral information both sequentially and retrospectively, thus creating a richer feature representation.
@@ III-A1 Bidirectional Processing Framework
In the SS non-linear Model, hyperspectral data $x(t)$ at each spectral band $t$, is projected into a hidden state representation $h(t)$ that integrates both forward and backward information. The forward and backward state transitions are modeled as follows:
@@
This bidirectional processing not only captures sequential dependencies, but also improves the discriminative power of the extracted features. By integrating forward and backward information, the SS non-linear Model effectively encapsulates the spectral structure in hyperspectral data, which is critical for accurate classification.
@@ III-B Proposed Method Overview and Architecture Description
The proposed SS non-linear Model architecture (illustrated in Figure 1) is designed to efficiently handle the high-dimensional data of hyperspectral images by dividing the processing into distinct blocks that target spectral and spatial information separately. This architecture consists of four main components: 1. Input Patch Preparation; 2. Hyperspectral Bi-Networks Block (for bidirectional spectral processing); 3. Spatial Feature Processing Block; 4. Classifier Block.
@@
Figure 1: The architectural overview of the Proposed SS non-linear Model model. The framework consists of four main components: (A) A hyperspectral image patch with dimensions $p\times p\times\text{CH}$; (B) The Hyperspectral Bi-networks Block; (C) The Spatial Processing Block; (D) The classifier. The process begins by extracting patches that serve as input to the SS non-linear Model block. This block includes a spatial processing stage that precedes the unique forward and backward operations, and the concatenation between the Bi-networks output and the spatial feature process is the input of the classifier block.
@@ III-B2 Hyperspectral BiNetwork Block
The hyperspectral bi-network block is central to the architecture of the SS non-linear Model. It processes the spectral information bidirectionally, utilizing forward and backward transformations to capture spectral dependencies in both directions. The following steps outline the processing within the Bi-Networks Block:
@@ III-B4 Classifier Block
III-B5 Algorithm for SS non-linear Model Operation
The SS non-linear Model efficiently processes hyperspectral image patches by integrating bidirectional spectral feature extraction and spatial processing to achieve accurate classification. The main steps in this operation involve Input Normalization, Forward and Backward Projections,Bidirectional Spectral Processing,Spatial Feature Processing, and Feature Concatenation and Classification.
Algorithm 1 SS non-linear Model Operation
@@ III-B5 Algorithm for SS non-linear Model Operation
5:Project $x_{\text{norm}}$ to $x_{\text{proj}}$ and $z_{\text{proj}}$ using non-linear transformations
@@ III-C Architectural Efficiency Analysis
The SS non-linear Model is designed to reduce computational demands while maintaining or improving classification accuracy. This section provides a quantitative analysis of the model’s efficiency, highlighting its advantages in terms of FLOPs, runtime, and overall computational complexity.
@@ III-C1 Floating Point Operations (FLOPs) Reduction
FLOPs provide a measure of the computational load associated with performing a single forward pass in a model. In hyperspectral image classification, conventional CNNs and transformers often involve high FLOPs due to extensive convolutional or self-attention operations across multiple spectral bands and spatial dimensions. The SS non-linear Model achieves a substantial reduction in FLOPs by adopting a bidirectional spectral processing approach without extensive kernel operations. Instead of using large convolutional filters or complex attention mechanisms, the SS non-linear Model leverages simpler 1D convolutions along with the efficient transformation matrices $A$ and $B$ resulting in lower computational complexity. To summarize, for a hyperspectral image with dimensions $Batch\times H\times W\times C$, where:
@@
The SS non-linear Model’s bidirectional approach enables it to reduce FLOPs by approximately 40% compared to conventional CNNs, assuming a kernel size $k>3$. This reduction is even more pronounced when compared to transformers, where the self-attention mechanism has quadratic complexity in the spectral dimension $CH$,
@@ III-C2 Runtime Efficiency
In practical runtime evaluations, the SS non-linear Model demonstrates a significant improvement in inference time over transformer-based architectures. This efficiency is primarily attributed to two architectural choices:
@@
Reduced Memory Footprint: The reduced number of parameters in the SS non-linear Model lowers memory usage, allowing it to perform faster, particularly on GPU and edge devices with limited memory resources.
When tested under identical hardware conditions, the SS non-Linear Model achieved a 30% faster inference time compared to transformer-based models, demonstrating its practical advantage for real-time applications.
@@ III-C3 Complexity Reduction
The bidirectional processing framework used in the SS non-Linear Model further reduces computational complexity by consolidating forward and backward spectral processing into a single framework. This approach contrasts with traditional models that handle spectral and spatial processing in isolation, often leading to redundant computations.
The SS non-Linear Model achieves efficiency through the following approximations:
@@
    SS non-Linear Model: Parameters: $O(CH+HW)$, FLOPs: $O(Batch.H.W\cdot CH)$.
The significantly lower FLOPs and parameter count in the SS non-Linear Model contribute to a leaner architecture, suitable for deployment in scenarios where computational resources are constrained.
@@ IV Experiments
This section presents a comprehensive evaluation of the proposed SS non-Linear Model, including dataset descriptions, experimental setup, benchmark comparisons, and in-depth analysis of classification performance. Our experiments examine model efficacy across several hyperspectral datasets, benchmarking SS non-Linear Model against state-of-the-art methods, and include ablation studies to validate its structural advantages.
@@ IV-A Datasets Description
Our experiments are conducted on three prominent hyperspectral datasets—Houston 2013, Indian Pines, and University of Pavia—each offering unique characteristics that make them ideal for evaluating the SS non-Linear Model’s performance across varied spectral and spatial complexities. The diversity of these datasets provides a comprehensive test of the model’s adaptability and robustness in hyperspectral image classification.
@@ 9 Shadows 103 923 1026
Each dataset’s specific spectral and spatial properties enable a well-rounded evaluation of the SS non-Linear Model, examining its robustness across complex, heterogeneous, and mixed-land-cover environments.
@@ IV-B Experimental Setup
To comprehensively assess the SS non-Linear Model’s classification effectiveness, we designed a robust experimental setup, encompassing evaluation metrics, benchmark comparisons, and specific implementation details. This setup provides a clear framework for evaluating both the accuracy and computational efficiency of the proposed model.
@@ IV-B1 Evaluation Metrics
The performance of the SS non-Linear Model is quantitatively assessed using a suite of standard metrics. Specifically, we employ Overall Accuracy ($OA$), which measures the general precision of the model; and the Kappa coefficient ($\kappa$), a statistical measure that accounts for chance agreement in classification tasks. Together, these metrics enable a robust evaluation of model performance.
@@ IV-B2 Benchmark Comparisons
We benchmark the SS non-Linear Model against a set of widely recognized hyperspectral image classification models, covering various deep learning architectures. These benchmarks encompass:
@@
To ensure consistency, comparative experiments are conducted with a standardized patch size of 15 and a 10% training data ratio. Additionally, we evaluate SS non-Linear Model with a smaller patch size of 3 to demonstrate its ability to maintain high accuracy with reduced computational resources, a key advantage in resource-limited environments.
@@ IV-B3 Implementation Details
The SS non-Linear Model is implemented using the PyTorch platform and trained on the school server based on GPU for enhanced computational performance. Optimization and Training: We use the Adam optimizer with a mini-batch size of 32 and a learning rate of 5e-4, with training set to 100 epochs across all datasets to balance training efficiency and model performance. Loss Function: Cross-entropy loss is chosen due to its effectiveness in multi-class classification tasks, particularly for hyperspectral data where class imbalances often require robust loss formulations. Data Augmentation: To improve model robustness, we apply geometric transformations, including rotations at 45°, 90°, and 135°, as well as vertical and horizontal flips, to the training data. This augmentation strategy enriches the training dataset by introducing additional spatial variability, enhancing the model’s ability to generalize across different spectral-spatial patterns. This experimental setup ensures that the SS non-Linear Model is rigorously tested under conditions that reflect both practical application scenarios and optimal evaluation standards.